# Omni-Video: Democratizing Unified Video Understanding and Generation

### v3

@@ 1 Introduction
The remarkable advancement of artificial intelligence in recent years, especially multimodal understanding chen2024internvl ; bai2025qwen25_vl empowered by large-language models (LLMs) and multimodal generationdalle3 ; sdxl ; wan2025wan models driven by diffusion models, has enabled thrilling possibilities to go beyond existing experiences across various modalities, including vision redmon2016you ; rombach2022high ; he2016deep ; dosovitskiyimage , language devlin2019bert ; brown2020language ; raffel2020exploring , as well as audio baevski2020wav2vec ; van2016wavenet ; park2019specaugment ; radford2023robust , etc. Despite these progresses, most understanding and generation systems are often designed separately, i.e., each specialized for a specific task or modality, leading to significant modeling redundancy and laborious computing complexity. Alternatively, unified multimodal models chen2025janus ; tong2024metamorph have emerged as a promising solution and gained increasing attention for their ability to input/output diverse data types seamlessly. Such unified pipeline is trained jointly across multiple modalities to model task-agnostic representations wang2025diffad ; chou2018unifying within a single framework, thereby removing tailored modality-specific components while achieving strong performance on both individual and compositional tasks hu2021unit ; wang2024git .
@@
Following this philosophy, recent efforts chen2025blip3 ; deng2025bagel have focused on developing unified architectures that simultaneously handle multimodal understanding and generation tasks. Considering that the transformer-based autoregressive (AR) architecture serves as the predominant architectural paradigm for multimodal understanding , earlier approaches project both texts and images into discrete tokens to facilitate AR modeling across modalities chen2025janus ; wang2024emu3 . For instance, Emu3 wang2024emu3 tokenize images, texts, and videos into a discrete space and train a single transformer model via next-token prediction on these multimodal sequences. Chameleon team2024chameleon designed an early fusion scheme to project all modalities into a share discrete representational space for better interaction and reasoning. Motivated by the great success of diffusion models sdxl ; dalle3 in visual generation, the most recent works chen2025blip3 ; deng2025bagel design hybrid diffusion-AR architectures that incorporate diffusion decoders to translate high-level visual tokens into photorealistic outputs. Representative works such as Transfusion zhou2024transfusion and Show-O xie2024show respectively employ continuous and discrete diffusion modeling with full attention for improved visual quality.
@@
However, training the above unified models from scratch requires substantial computational resources and a massive amounts of multimodal data. Additionally, integrating MLLMs and diffusion models within a unified system poses significant challenges in architectural modifications and hyperparameter tuning. To tackle these challenges, MetaMorph tong2024metamorph demonstrated that an improved visual understanding inherently endows MLLMs with visual generation capabilities. Building on this insight, a straightforward visual predictive instruction tuning (VPiT) strategy was introduced to train LLMs to generate continuous visual tokens, which are subsequently decoded into plausible images using a modified diffusion model. Accordingly, MetaMorph achieved competitive performance on both image understanding and generation tasks with minimal architectural changes and computational overhead. Nevertheless, previous unified models have predominantly targeted static images and show limited capability when it comes to processing videos, where accurately modeling temporal dynamics, ensuring motion coherence, and maintaining frame-wise consistency are essential. Furthermore, the elevated data and computational demands for video modeling impose significant scalability constraints on current unified frameworks. Consequently, delivering an appropriate framework for unified video understanding and generation is critically important for immersive media understanding, production, and even interaction.
@@
To fill the vacant, this technical report proposes Omni-Video, an efficient and effective framework for unified video understanding and generation that combines a lightweight architectural design with a multi-stage training scheme. Borrowing the exceptional video understanding and generation capabilities of existing MLLMs (e.g., VILA lin2024vila ) and text-to-video (T2V) diffusion (e.g., Wan wan2025wan ) models, our method aims at teaching existing video understanding MLLMs to produce video tokens that are ‘understandable’ for existing T2V diffusion models. More specifically, we first tuning MLLMs to produce textual tokens and visual tokens with separate text and vision heads following tong2024metamorph , the vision head generates continuous video tokens that is aligned with the representation of a pretrained vision encoder, e.g., SigLIP2 tschannen2025siglip2 . Then, we adapt the continuous video tokens into the T2V diffusion model’s conditional space with a dedicated lightweight adapter, such that the T2V model could synthesize faithful videos directly from the continuous video tokens. Notably, only a small subset of model parameters, i.e., the text and vision head in stage-1 and the adapter in stage-2, are optimized. To further activate the potential visual generation capability of MLLMs, we unlock the parameters of text and vision head, adapter, as well as the T2V model and train the whole system with meticulously curated high-quality understanding and generation data, thereby improving the synthesis quality and the understanding ability. This selective fine-tuning strategy requires only minimal data and computational resources, allowing for efficient adaptation without compromising the foundational performance of the pretrained model.
@@
In addition to generate harmonious videos conditioned on user instructions, we also extend our unified framework towards detail-preserving visual editing scenarios. One straightforward solution is to feed original signals into the visual encoder of MLLMs and retrain the unified model to reconstruct unedited areas and penalize editing areas with paired editing datayu2024promptfix ; zhao2024ultraedit ; zhang2024hive . This approach, however, incurs substantial computational overhead with massive data and often leads to suboptimal performance as MLLMs primary focus on high-level semantic features rather low-level details. Since T2V diffusion models are trained to recover clean videos from corrupted signals, they inherently excel at processing fine-grained visual information. Accordingly, we leverage pretrained VAEs to extract low-level compressed priors from original input and directly send them to the diffusion model. Such that, the original signals interact with noised embeddings throughout all denoising steps, ensuring high-fidelity visual preservation during the editing process. Furthermore, motivated by the recent chain-of-thought reasoning capabilities in the LLMs community, we devise a “think mode" to achieve higher-quality visual outputs through logical inference and contextual planning in video generation and editing. Together, our framework enables seamlessly predicting multimodal tokens for video understanding, generation, as well as video editing tasks, with minimal architectural modification, diverse data sources, and computation-efficient optimization.
@@ 2.1 Multimodal Understanding
Taking advantage of advanced large language models (LLMs) achiam2023gpt4 ; grattafiori2024llama3 ; yang2025qwen3 , multimodal large language models (MLLMs) have gained extensive research attention and demonstrated magnificent capabilities in understanding and reasoning multimodal content li2024llava-ov ; bai2025qwen25_vl ; guo2025seed15_vl . In order to inject knowledge beyond the text modality into LLMs, existing methods typically employ modality-specific encoders (e.g., CLIP radford2021clip ) to extract latent embeddings and perform alignment within the LMMs’ transformer space. For instance, LLAVA liu2023llava ; liu2024llava_1_5 and MiniGPT-4 zhu2023minigpt4 connect pre-trained visual encoders and LLMs with simple projection layers, enabling LLMs to perform complex multimodal understanding while preserving their textual reasoning ability. Building this insights, more advanced techniques such as scaling the visual encoders chen2024internvl , support dynamic and high input resolutions wang2024qwen2-vl , employment of Mixture-of-Experts (MOE) architectures jiang2024mixtral ; liu2024deepseekv2 , replacing linear projection layers with different architectures li2023blip2 ; alayrac2022flamingo , improvement of training receipts karamcheti2024prismatic ; lin2024vila are developed to enhance the fine-grained perception capabilities of MLLMs. Similarly, video-centric MLLMs lin2024vila encode videos into sequential visual tokens that are compatible with LLMs. Although these approaches excel at multimodal understanding and reasoning, they primarily focus on perception tasks and lack the capability of producing other modalities beyond text.
@@ 2.2 Multimodal Generation
Early multimodal generation models employ generative adversarial networks (GANs) goodfellow2014gan to synthesis novel outputs. However, the scalability, visual quality and resolution of GANs are relatively poor. Diffusion models (DMs) DDPM and autoregressive (AR) models box1970time ; radford2018gpt , equipped with their inherent stable optimization and scalability, the transformer architecture, large-scale training and infrastructure, and trillion-level multimodal corpus, are current state-of-the-art for image/video synthesis.
@@
Diffusion models (DMs). DMs gradually transform simple noise distributions into complex data distributions through a series of learned denoising steps. The pioneering work laid the theoretical foundations, while numerous subsequent studies have enhanced the efficiency and output quality through improved training/sampling strategies lu2022dpm ; karras2024analyzing ; sdxl ; gao2025seedream and optimization techniques lu2024simplifying ; song2023improved ; liu2024playground , etc. In particular, the introduction of stable diffusion (SD) and diffusion transformer (DiT)  peebles2023dit directly caused the blooming of high-quality, high-resolution, and long-duration visual generation. For instance, Sora openai2024sora , built on the diffusion transformer architecture, achieved superior performance in producing detailed, high-resolution videos that adhere to the user’s instructions. Follow-up approaches including open-sourcedzheng2024opensora ; tan2025raccoon ; wan2025wan and closed-source models klingai2024 ; gao2025seedance have further advanced the state-of-the-art in text-to-video generation. These developments underscore that enhanced diffusion techniques are enabling richer, more dynamic, and semantically coherent visual content creation. In this work, we integrate the open-sourced Wan2.1 wan2025wan video generation model into our unified understanding and generation system as its leading performance in the open-source community.
@@
Autoregressive (AR) models. Borrowing the success of AR modeling in language domain, another line of approaches learn to synthesize visual content in autoregressive manner. Concretely, visual inputs are first mapped into a sequence of discrete tokens, after which both visual and text tokens are modeled via next token prediction. To encode continuous visual signals into discrete tokens, vector quantization techniquesyu2023magvitv2 ; tian2024var ; li2024mar are usually applied to high-dimensional data into a finite set of latent codes (codebook). DALLE dalle2 ; dalle3 demonstrated that scaling with sufficient data autoregressively could generate high-quality images from textual descriptions. MAGVIT-v2 yu2023magvitv2 showed that GPT-style LLMs might outperform diffusion models on visual generation benchmarks with an improved video tokenizer. With this top-performing tokenizer, VideoPoet further identified the potential of AR model, which was trained on discrete visual, text and audio latents, in producing plausible videos, especially excel in modeling complex large motions. Differently, VAR tian2024var designed a coarse-to-fine “next-scale prediction" paradigm to learn visual distributions and MAR li2024mar applied AR modeling in the continuous space by modeling per-token distribution by diffusion.
@@ 2.3 Unified Multimodal Understanding and Generation
In an effort to build a single system for both understanding and generation tasks, substantial developments have been witnessed in unified multimodal models recently. In addition to the text modality, these models process various modalities as input (e.g., image, video) and also output various modalities in a unified manner. Depending on whether visual signals are encoded into discrete or continuous tokens, existing unified multimodal models primarily consist of pure AR models chen2025janus ; wang2024emu3 ; wu2025liquid and diffusion-AR hybrids tong2024metamorph ; chen2025blip3 ; liao2025mogao . For example, both Chameleon team2024chameleon and Emu3 wang2024emu3 represented image/videos with VQ-VAE van2017vqvae as discrete tokens and trained uniform transformer-based LLM models from scratch in next token prediction manner. Janus and Janus-pro chen2025janus decoupled visual encoding with SigLIP zhai2023siglip and VQ-VAE sun2024llamagen for multimodal understanding and generation, respectively. Such decoupling ameliorated the conflict of visual encoder for understanding and generation tasks. Following this philosophy, the most recent work converted images into discrete tokens with a text-aligned tokenizer that was aligned with LLM embeddings. However, discreting visual signals often struggles to preserve fine-grained details, leading to sub-optimal synthesis performance. Consequently, UniFluid fan2025unifluid build a unified framework from continuous visual tokens following the MAR’s per-token distribution modeling.
Alternatively, other approaches integrate diffusion models into LLMs for visual decoding. In particular, Show-O xie2024show and Tranfusion zhou2024transfusion unified diffusion models for generation tasks and LLMs for textual responses, while the former adopted discrete diffusion and the latter used continuous diffusion. Along this line, Show-O2 xie2025show2 and Mogao liao2025mogao improved the model performance via combining AR modeling and flow-matching lipman2022flow . In order to identify the impact of core components (i.e., image representation, training objective, training strategies) of hybrid diffusion-AR unified models, BLIP3-o chen2025blip3 presented a thorough empirical study. Despite these models’ success in unifying understanding and generation, training these models from scratch requires massive compute and data resources. The most related work to ours is MetaMorph, which trained LLMs to predict multimodal tokens for generation with visual-predictive instruction tuning (VPiT). However, most of existing methods, including MetaMorph, focus on modeling static images and show very limited ability for processing videos. Therefore, this work presents an efficient and effective framework for unified video understanding and generation.

### v4

@@ 1 Introduction
The remarkable advancement of artificial intelligence in recent years, especially multimodal understanding Chen et al. (2024c); Bai et al. (2025) empowered by large-language models (LLMs) and multimodal generationBetker et al. (2023); Podell et al. (2023); Wan et al. (2025) models driven by diffusion models, has enabled thrilling possibilities to go beyond existing experiences across various modalities, including vision Redmon et al. (2016); Rombach et al. (2022); He et al. (2016); Dosovitskiy et al. , language Devlin et al. (2019); Brown et al. (2020); Raffel et al. (2020), as well as audio Baevski et al. (2020); Dieleman et al. (2016); Park et al. (2019); Radford et al. (2023), etc. Despite these progresses, most understanding and generation systems are often designed separately, i.e., each specialized for a specific task or modality, leading to significant modeling redundancy and laborious computing complexity. Alternatively, unified multimodal models Chen et al. (2025b); Tong et al. (2024b) have emerged as a promising solution and gained increasing attention for their ability to input/output diverse data types seamlessly. Such unified pipeline is trained jointly across multiple modalities to model task-agnostic representations Wang et al. (2025); Chou et al. (2018) within a single framework, thereby removing tailored modality-specific components while achieving strong performance on both individual and compositional tasks Hu and Singh (2021); Wang et al. (2024a).
@@
Following this philosophy, recent efforts Chen et al. (2025a); Deng et al. (2025) have focused on developing unified architectures that simultaneously handle multimodal understanding and generation tasks. Considering that the transformer-based autoregressive (AR) architecture serves as the predominant architectural paradigm for multimodal understanding , earlier approaches project both texts and images into discrete tokens to facilitate AR modeling across modalities Chen et al. (2025b); Wang et al. (2024c). For instance, Emu3 Wang et al. (2024c) tokenize images, texts, and videos into a discrete space and train a single transformer model via next-token prediction on these multimodal sequences. Chameleon Team (2024) designed an early fusion scheme to project all modalities into a share discrete representational space for better interaction and reasoning. Motivated by the great success of diffusion models Podell et al. (2023); Betker et al. (2023) in visual generation, the most recent works Chen et al. (2025a); Deng et al. (2025) design hybrid diffusion-AR architectures that incorporate diffusion decoders to translate high-level visual tokens into photorealistic outputs. Representative works such as Transfusion Zhou et al. (2024) and Show-O Xie et al. (2024) respectively employ continuous and discrete diffusion modeling with full attention for improved visual quality.
@@
However, training the above unified models from scratch requires substantial computational resources and a massive amounts of multimodal data. Additionally, integrating MLLMs and diffusion models within a unified system poses significant challenges in architectural modifications and hyperparameter tuning. To tackle these challenges, MetaMorph Tong et al. (2024b) demonstrated that an improved visual understanding inherently endows MLLMs with visual generation capabilities. Building on this insight, a straightforward visual predictive instruction tuning (VPiT) strategy was introduced to train LLMs to generate continuous visual tokens, which are subsequently decoded into plausible images using a modified diffusion model. Accordingly, MetaMorph achieved competitive performance on both image understanding and generation tasks with minimal architectural changes and computational overhead. Nevertheless, previous unified models have predominantly targeted static images and show limited capability when it comes to processing videos, where accurately modeling temporal dynamics, ensuring motion coherence, and maintaining frame-wise consistency are essential. Furthermore, the elevated data and computational demands for video modeling impose significant scalability constraints on current unified frameworks. Consequently, delivering an appropriate framework for unified video understanding and generation is critically important for immersive media understanding, production, and even interaction.
@@
To fill the vacant, this technical report proposes Omni-Video, an efficient and effective framework for unified video understanding and generation that combines a lightweight architectural design with a multi-stage training scheme. Borrowing the exceptional video understanding and generation capabilities of existing MLLMs (e.g., VILA Lin et al. (2024)) and text-to-video (T2V) diffusion (e.g., Wan Wan et al. (2025)) models, our method aims at teaching existing video understanding MLLMs to produce video tokens that are ‘understandable’ for existing T2V diffusion models. More specifically, we first tuning MLLMs to produce textual tokens and visual tokens with separate text and vision heads following Tong et al. (2024b), the vision head generates continuous video tokens that is aligned with the representation of a pretrained vision encoder, e.g., SigLIP2 Tschannen et al. (2025). Then, we adapt the continuous video tokens into the T2V diffusion model’s conditional space with a dedicated lightweight adapter, such that the T2V model could synthesize faithful videos directly from the continuous video tokens. Notably, only a small subset of model parameters, i.e., the text and vision head in stage-1 and the adapter in stage-2, are optimized. To further activate the potential visual generation capability of MLLMs, we unlock the parameters of text and vision head, adapter, as well as the T2V model and train the whole system with meticulously curated high-quality understanding and generation data, thereby improving the synthesis quality and the understanding ability. This selective fine-tuning strategy requires only minimal data and computational resources, allowing for efficient adaptation without compromising the foundational performance of the pretrained model.
@@
In addition to generate harmonious videos conditioned on user instructions, we also extend our unified framework towards detail-preserving visual editing scenarios. One straightforward solution is to feed original signals into the visual encoder of MLLMs and retrain the unified model to reconstruct unedited areas and penalize editing areas with paired editing dataYu et al. (2024b); Zhao et al. (2024); Zhang et al. (2024a). This approach, however, incurs substantial computational overhead with massive data and often leads to suboptimal performance as MLLMs primary focus on high-level semantic features rather low-level details. Since T2V diffusion models are trained to recover clean videos from corrupted signals, they inherently excel at processing fine-grained visual information. Accordingly, we leverage pretrained VAEs to extract low-level compressed priors from original input and directly send them to the diffusion model. Such that, the original signals interact with noised embeddings throughout all denoising steps, ensuring high-fidelity visual preservation during the editing process. Furthermore, motivated by the recent chain-of-thought reasoning capabilities in the LLMs community, we devise a “think mode" to achieve higher-quality visual outputs through logical inference and contextual planning in video generation and editing. Together, our framework enables seamlessly predicting multimodal tokens for video understanding, generation, as well as video editing tasks, with minimal architectural modification, diverse data sources, and computation-efficient optimization.
@@ 2.1 Multimodal Understanding
Taking advantage of advanced large language models (LLMs) Achiam et al. (2023); Grattafiori et al. (2024); Yang et al. (2025), multimodal large language models (MLLMs) have gained extensive research attention and demonstrated magnificent capabilities in understanding and reasoning multimodal content Li et al. (2024a); Bai et al. (2025); Guo et al. (2025). In order to inject knowledge beyond the text modality into LLMs, existing methods typically employ modality-specific encoders (e.g., CLIP Radford et al. (2021)) to extract latent embeddings and perform alignment within the LMMs’ transformer space. For instance, LLAVA Liu et al. (2023, 2024d) and MiniGPT-4 Zhu et al. (2023a) connect pre-trained visual encoders and LLMs with simple projection layers, enabling LLMs to perform complex multimodal understanding while preserving their textual reasoning ability. Building this insights, more advanced techniques such as scaling the visual encoders Chen et al. (2024c), support dynamic and high input resolutions Wang et al. (2024b), employment of Mixture-of-Experts (MOE) architectures Jiang et al. (2024); Liu et al. (2024a), replacing linear projection layers with different architectures Li et al. (2023); Alayrac et al. (2022), improvement of training receipts Karamcheti et al. (2024); Lin et al. (2024) are developed to enhance the fine-grained perception capabilities of MLLMs. Similarly, video-centric MLLMs Lin et al. (2024) encode videos into sequential visual tokens that are compatible with LLMs. Although these approaches excel at multimodal understanding and reasoning, they primarily focus on perception tasks and lack the capability of producing other modalities beyond text.
@@ 2.2 Multimodal Generation
Early multimodal generation models employ generative adversarial networks (GANs) Goodfellow et al. (2014) to synthesis novel outputs. However, the scalability, visual quality and resolution of GANs are relatively poor. Diffusion models (DMs) Ho et al. (2020) and autoregressive (AR) models Box et al. (1970); Radford et al. (2018), equipped with their inherent stable optimization and scalability, the transformer architecture, large-scale training and infrastructure, and trillion-level multimodal corpus, are current state-of-the-art for image/video synthesis.
@@
Diffusion models (DMs). DMs gradually transform simple noise distributions into complex data distributions through a series of learned denoising steps. The pioneering work laid the theoretical foundations, while numerous subsequent studies have enhanced the efficiency and output quality through improved training/sampling strategies Lu et al. (2022); Karras et al. (2024); Podell et al. (2023); Gao et al. (2025a) and optimization techniques Lu and Song (2024); Song and Dhariwal (2023); Liu et al. (2024b), etc. In particular, the introduction of stable diffusion (SD) and diffusion transformer (DiT)  Peebles and Xie (2023) directly caused the blooming of high-quality, high-resolution, and long-duration visual generation. For instance, Sora OpenAI (2024), built on the diffusion transformer architecture, achieved superior performance in producing detailed, high-resolution videos that adhere to the user’s instructions. Follow-up approaches including open-sourcedZheng et al. (2024); Tan et al. (2025); Wan et al. (2025) and closed-source models AI (2024); Gao et al. (2025b) have further advanced the state-of-the-art in text-to-video generation. These developments underscore that enhanced diffusion techniques are enabling richer, more dynamic, and semantically coherent visual content creation. In this work, we integrate the open-sourced Wan2.1 Wan et al. (2025) video generation model into our unified understanding and generation system as its leading performance in the open-source community.
@@
Autoregressive (AR) models. Borrowing the success of AR modeling in language domain, another line of approaches learn to synthesize visual content in autoregressive manner. Concretely, visual inputs are first mapped into a sequence of discrete tokens, after which both visual and text tokens are modeled via next token prediction. To encode continuous visual signals into discrete tokens, vector quantization techniquesYu et al. (2023); Tian et al. (2024); Li et al. (2024b) are usually applied to high-dimensional data into a finite set of latent codes (codebook). DALLE Ramesh et al. (2022); Betker et al. (2023) demonstrated that scaling with sufficient data autoregressively could generate high-quality images from textual descriptions. MAGVIT-v2 Yu et al. (2023) showed that GPT-style LLMs might outperform diffusion models on visual generation benchmarks with an improved video tokenizer. With this top-performing tokenizer, VideoPoet further identified the potential of AR model, which was trained on discrete visual, text and audio latents, in producing plausible videos, especially excel in modeling complex large motions. Differently, VAR Tian et al. (2024) designed a coarse-to-fine “next-scale prediction" paradigm to learn visual distributions and MAR Li et al. (2024b) applied AR modeling in the continuous space by modeling per-token distribution by diffusion.
@@ 2.3 Unified Multimodal Understanding and Generation
In an effort to build a single system for both understanding and generation tasks, substantial developments have been witnessed in unified multimodal models recently. In addition to the text modality, these models process various modalities as input (e.g., image, video) and also output various modalities in a unified manner. Depending on whether visual signals are encoded into discrete or continuous tokens, existing unified multimodal models primarily consist of pure AR models Chen et al. (2025b); Wang et al. (2024c); Wu et al. (2025) and diffusion-AR hybrids Tong et al. (2024b); Chen et al. (2025a); Liao et al. (2025). For example, both Chameleon Team (2024) and Emu3 Wang et al. (2024c) represented image/videos with VQ-VAE Van Den Oord et al. (2017) as discrete tokens and trained uniform transformer-based LLM models from scratch in next token prediction manner. Janus and Janus-pro Chen et al. (2025b) decoupled visual encoding with SigLIP Zhai et al. (2023) and VQ-VAE Sun et al. (2024) for multimodal understanding and generation, respectively. Such decoupling ameliorated the conflict of visual encoder for understanding and generation tasks. Following this philosophy, the most recent work converted images into discrete tokens with a text-aligned tokenizer that was aligned with LLM embeddings. However, discreting visual signals often struggles to preserve fine-grained details, leading to sub-optimal synthesis performance. Consequently, UniFluid Fan et al. (2025a) build a unified framework from continuous visual tokens following the MAR’s per-token distribution modeling.
Alternatively, other approaches integrate diffusion models into LLMs for visual decoding. In particular, Show-O Xie et al. (2024) and Tranfusion Zhou et al. (2024) unified diffusion models for generation tasks and LLMs for textual responses, while the former adopted discrete diffusion and the latter used continuous diffusion. Along this line, Show-O2 Xie et al. (2025) and Mogao Liao et al. (2025) improved the model performance via combining AR modeling and flow-matching Lipman et al. (2022). In order to identify the impact of core components (i.e., image representation, training objective, training strategies) of hybrid diffusion-AR unified models, BLIP3-o Chen et al. (2025a) presented a thorough empirical study. Despite these models’ success in unifying understanding and generation, training these models from scratch requires massive compute and data resources. The most related work to ours is MetaMorph, which trained LLMs to predict multimodal tokens for generation with visual-predictive instruction tuning (VPiT). However, most of existing methods, including MetaMorph, focus on modeling static images and show very limited ability for processing videos. Therefore, this work presents an efficient and effective framework for unified video understanding and generation.